### v1

# Eigenvalue Based Active User Enumeration for Grant-Free Access Under Carrier Frequency Offsets

Abstract
This paper investigates a grant-free non-orthogonal multiple access (GF-NOMA) system in the presence of carrier frequency offsets. We propose two schemes for enumerating active users in such a GF-NOMA system, which is equivalent to estimating the sparsity level. Both schemes utilize a short common pilot and the eigenvalues of the sample covariance matrix of the received signal. The two schemes differ in their treatment of noise variance: one exploits known variance information, while the other is designed to function without this knowledge. Simulation results demonstrate the effectiveness of the proposed schemes in terms of the normalized mean-squared error.
Index Terms:
Grant-free access, active user enumeration, sparsity level estimation, carrier frequency offsets.
I Introduction
Future wireless communication systems are required to support time-sensitive applications, such as robot control and automated driving, which involve sporadic data traffic due to massive short-packet communications[1, 2, 3]. The conventional multiple access schemes employing grant-based approaches are unable to support such scenarios due to the non-negligible delay induced by the handshake procedure for grant acquisition[4].
As a promising solution to this problem, grant-free random access schemes such as grant-free non-orthogonal multiple access (GF-NOMA) in which the base station (BS) does not exclusively assign radio resources to active users for data transmission have attracted significant attention [4, 5]. In a GF-NOMA system, each active user can directly transmit its packet to the BS without waiting for permission; as a result, however, the BS must cope with active user detection (AUD) to recovery transmitted data accurately.
In this context, many approaches to accurate AUD have been investigated for synchronous GF-NOMA systems, where perfect synchronization between the BS and users in both time and frequency is assumed[6, 7, 8, 9, 10, 11, 12, 13, 14]. On the other hand, achieving ideal synchronization is impractical due to the nature of grant-free transmission, and AUD for asynchronous scenarios remains an ongoing research area[4]. In practice, due to the practical limitation of the crystal oscillators[15], the resulting carrier frequency offsets (CFOs) among the users degrade the performance.
To overcome this impediment, some approaches to AUD in the presence of CFOs have recently been proposed[16, 17, 18, 19, 20]. For instance, [16] proposed the block coordinate descent (BCD)-based method for AUD and channel estimation (CE). In addition, [17] and [18] handled AUD using maximum likelihood estimation (MLE), which achieves higher accuracy than the BCD-based method of [16]. Meanwhile, since the MLE-based method use a discrete approximation of the rotation caused by CFOs, the improvement of the performance requires the increase in the discretization size, resulting in high computational complexity. As an alternative, the approximate message passing (AMP)-based approaches for AUD and CE have been proposed in [19] and [20], assuming that the frequency error is limited to a certain range that meets standard requirements. Although they can perform AUD and CE efficiently, prior information such as activity ratio is necessary. This implies that a scheme to estimate prior information is required to achieve efficient AUD in the presence of CFOs.
For synchronous scenarios, schemes to estimate activity ratio, namely the number of active users, have been proposed[9, 11, 14]. Throughout this paper, this process is referred to as active user enumeration (AUE). In [9], the number of active users is estimated through rank estimation of the covariance matrix of the received pilot signal, to achieve precise AUD using the Riemannian optimization. However, this approach necessitates the eigenvalue decomposition of the received pilot signal and a sufficiently long pilot sequence, resulting in high computational complexity for AUE. On the other hand, [11] and [14] have proposed the AUE schemes using a very short common pilot sequence. The scheme in [11] uses a sequence orthogonal to the common sequence, while the scheme in [14] considers the MLE problem of the number of active users based on the covariance matrix of the received signal. These AUE schemes, utilizing a short common pilot, can estimate the number of active users with low complexity.
To this end, in this letter, we propose two AUE schemes using a short common pilot sequence for estimating the number of active users in the presence of CFOs. Specifically, these AUE schemes that use the eigenvalues of the sample covariance matrix of the received signal, with one scheme designed for scenarios where the noise variance information is known, and the other for when it is unknown. Numerical results demonstrate that the proposed AUE schemes achieve superior estimation accuracy compared to the conventional schemes proposed in [11] and [14].
Notation: Throughout this paper, we use the following notation. The transpose and conjugate transpose are denoted by $(\cdot)^{\mathrm{T}}$ and $(\cdot)^{\mathrm{H}}$, respectively. The $l_{p}$-norm of a vector is expressed as $\|\cdot\|_{p}$. In addition, $\mathbf{0}_{L}$ and $\mathbf{I}_{L}$ denote the $L$-dimensional column vectors in which all elements are zeros and the $L\times L$ identity matrix, respectively. The determinant and trace of a square matrix $\mathbf{A}$ are denoted by $|\mathbf{A}|$ and $\mathrm{tr}\left(\mathbf{A}\right)$. The expectation operator and operation that rounds $x$ to the nearest integer are represented by $\mathbb{E}[\cdot]$ and $\lfloor x\rceil$, respectively. Finally, $\mathcal{CN}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ represents a multivariate complex Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$.
II System Model
[…]
Let $\mathbf{Y}\in\mathbb{C}^{2\times M}$ denote the received pilot signal at the BS. Then, the received signal can be written as
$\mathbf{Y}$ $=\sum_{n\in\mathcal{A}}\sqrt{\beta_{n}\rho_{n}}(\mathbf{s}\circ\boldsymbol{\tau}(\omega_{n}))\mathbf{h}^{\mathrm{T}}_{n}+\mathbf{Z}$
$=\sum_{n\in\mathcal{A}}\sqrt{\beta_{n}\rho_{n}}\boldsymbol{\tau}(\omega_{n})\mathbf{h}^{\mathrm{T}}_{n}+\mathbf{Z},$ (1)
where $\boldsymbol{\tau}(\omega_{n})=[1\ e^{j\omega_{n}}]^{\mathrm{T}}\in\mathbb{C}^{2\times 1}$ is the phase rotation vector due to $\omega_{n}$, and $\mathcal{A}\subset\{1,2,\ldots,N\}$, whose cardinality is $K$, is the set of active users. In addition, $\mathbf{h}_{n}\sim\mathcal{CN}(\mathbf{0}_{M},\mathbf{I}_{M}))$ and $\mathbf{Z}\in\mathbb{C}^{2\times M}$ denote the small-scale fading channel vector between user $n$ and BS, and the noise matrix whose each entry follows an independently and identically distributed (i.i.d.) complex Gaussian distribution with zero mean and variance $\sigma^{2}_{\mathrm{z}}$, respectively. $\rho_{n}$ and $\beta_{n}$ denote the transmit power of user $n$ and the large-scale fading component between user $n$ and the BS. Throughout this paper, each active user is assumed to compensate for the effect of the large-scale fading component by the uplink power control, i.e., $\beta_{n}\rho_{n}=1$[11]. Thus, the received pilot signal model can be simplified as
$\mathbf{Y}=\sum_{n\in\mathcal{A}}\boldsymbol{\tau}(\omega_{n})\mathbf{h}^{\mathrm{T}}_{n}+\mathbf{Z}.$ (2)
Our goal is to estimate the number of active users, $K$, from the received pilot signal taking the effect of CFOs into account.
III Conventional AUE Schemes
In this section, to clarify the difference between conventional and proposed AUE schemes, we review the AUE schemes proposed in [11] and [14]. These conventional schemes utilize the assumption that the probability function of $\mathbf{Y}$ given $K$ and $\mathbf{s}$ is given by
$p(\mathbf{Y}|K,\mathbf{s})=\frac{1}{|\pi\boldsymbol{\Sigma}|^{M}}\prod^{M}_{m=1}\exp\left(-\mathbf{y}^{\mathrm{H}}_{m}\boldsymbol{\Sigma}^{-1}\mathbf{y}^{\mathrm{H}}_{m}\right),$ (3)
where
$\boldsymbol{\Sigma}=\mathbb{E}\left[\mathbf{y}_{m}\mathbf{y}^{\mathrm{H}}_{m}\right]=K\mathbf{s}\mathbf{s}^{\mathrm{H}}+\sigma^{2}_{\mathrm{z}}\mathbf{I}_{2}.$ (4)
Note that this assumption does not consider the effect of CFOs.
The AUE scheme in [11] utilizes the sequence $\mathbf{s}_{\perp}\in\mathbb{C}^{2\times 1}$ that is orthogonal to $\mathbf{s}$ and satisfies $\|\mathbf{s}_{\perp}\|_{2}=\|\mathbf{s}\|_{2}$. Let $\mathbf{R}\triangleq\mathbf{Y}\mathbf{Y}^{\mathrm{H}}/M\in\mathbb{C}^{2\times 2}$ be the sample covariance matrix of the received pilot signal in (2). Then, the criterion to estimate $K$ in [11] can be expressed as
$\hat{K}_{\mathrm{orth}}=\Bigg{\lfloor}\frac{\mathbf{s}^{\mathrm{H}}\mathbf{R}\mathbf{s}}{\|\mathbf{s}\|^{4}_{2}}-\frac{\mathbf{s}^{\mathrm{H}}_{\perp}\mathbf{R}\mathbf{s}_{\perp}}{\|\mathbf{s}\|^{2}_{2}\|\mathbf{s}_{\perp}\|^{2}_{2}}\Bigg{\rceil}.$ (5)
When we set $\|\mathbf{s}\|_{2}=\|\mathbf{s}_{\perp}\|_{2}=1$, (5) is equivalent to [11, eq. (5)]. It is worth noting that this scheme does not require the knowledge of noise variance and that $\hat{K}_{\mathrm{orth}}$ converges to $K$ when $M\to\infty$ and $\omega_{n}=0$.
On the other hand, the AUE scheme in [14] considers the MLE problem of $K$. This problem can be formulated as
$\underset{K\in[0,N]}{\mathrm{minimize}}$ $\quad\log|\boldsymbol{\Sigma}|+\mathrm{tr}\left(\boldsymbol{\Sigma}^{-1}\mathbf{R}\right)$ (6)
The solution of the optimization problem can be obtained by
$\hat{K}_{\mathrm{MLE}}=\Bigg{\lfloor}\frac{\mathbf{s}^{\mathrm{H}}\mathbf{R}\mathbf{s}}{\|\mathbf{s}\|^{4}_{2}}-\frac{\sigma^{2}_{\mathrm{z}}}{\|\mathbf{s}\|^{2}_{2}}\Bigg{\rceil}.$ (7)
As is obvious from (7), the estimation of $K$ in [14] does not rely on the sequence orthogonal to $\mathbf{s}$, while it requires the knowledge of noise variance.
IV Proposed AUE Schemes
In this section, we propose two AUE schemes using the sample covariance matrix of the received signal. Specifically, the proposed schemes utilize the eigenvalues of $\mathbf{R}$. Hereafter, $\lambda_{\max}$ and $\lambda_{\min}\ (<\lambda_{\max})$ denote the eigenvalues.
IV-A Preliminaries
The proposed schemes consider the covariance matrix of $\mathbf{y}_{m}$ given $\omega_{n}$ for $n\in\mathcal{A}$, i.e.,
$\tilde{\boldsymbol{\Sigma}}=\mathbb{E}\left[\mathbf{y}_{m}\mathbf{y}^{\mathrm{H}}_{m}\right]$ $=\sum_{n\in\mathcal{A}}\boldsymbol{\tau}(\omega_{n})\boldsymbol{\tau}(\omega_{n})^{\mathrm{H}}+\sigma^{2}_{\mathrm{z}}\mathbf{I}_{2}$
$=\begin{bmatrix}K+\sigma^{2}_{\mathrm{z}}&\sum_{n\in\mathcal{A}}e^{-j\omega_{n}}\\ \sum_{n\in\mathcal{A}}e^{j\omega_{n}}&K+\sigma^{2}_{\mathrm{z}}\\ \end{bmatrix}$ (8)
The eigenvalues of this covariance matrix are given by
$\Lambda_{\max}$ $=K+\sigma^{2}_{\mathrm{z}}+\gamma,$ (9)
$\Lambda_{\min}$ $=K+\sigma^{2}_{\mathrm{z}}-\gamma,$ (10)
where
$\gamma=\sqrt{\left(\sum_{n\in\mathcal{A}}e^{j\omega_{n}}\right)\left(\sum_{n\in\mathcal{A}}e^{-j\omega_{n}}\right)}.$ (11)
Assuming $K$ is sufficiently large, $\gamma$ can be approximated as
$\gamma$ $=K\sqrt{\left(\frac{1}{K}\sum_{n\in\mathcal{A}}e^{j\omega_{n}}\right)\left(\frac{1}{K}\sum_{n\in\mathcal{A}}e^{-j\omega_{n}}\right)}$
$\approx K\sqrt{\mathbb{E}\left[e^{j\omega}\right]\mathbb{E}\left[e^{-j\omega}\right]},$ (12)
where $\omega$ represents the random variable following a uniform distribution in the interval $[-2\pi\epsilon_{\max},2\pi\epsilon_{\max}]$. $\mathbb{E}\left[e^{j\omega}\right]$ in (12) is the characteristic function of $\omega$, which is given by
[…]
$=\frac{\sin(2\pi\epsilon_{\max})}{2\pi\epsilon_{\max}}$
$=\mathrm{sinc}(2\pi\epsilon_{\max}).$ (13)
IV-B Proposed AUE Scheme With Knowledge of $\sigma^{2}_{\mathrm{z}}$
As is obvious from (9) and (10), $K$ can be ideally obtained based on the sum of eigenvalues of $\tilde{\boldsymbol{\Sigma}}$, i.e., $K=(\Lambda_{\max}+\Lambda_{\min})/2-\sigma^{2}_{\mathrm{z}}$, if the information of $\sigma^{2}_{\mathrm{z}}$ is available at the BS. Following this fact, one of the proposed schemes obtains the estimate of $K$ by
$\hat{K}_{\mathrm{prop1}}=\frac{\lambda_{\max}+\lambda_{\min}}{2}-\sigma^{2}_{\mathrm{z}}.$ (14)
Since this scheme is based on the calculation to offset the term including CFOs, $\gamma$, it is expected to be robust against CFOs.
IV-C Proposed AUE Scheme Without Knowledge of $\sigma^{2}_{\mathrm{z}}$
Since $\mathbb{E}[e^{j\omega}]$ is an even function, $\mathbb{E}[e^{-j\omega}]=\mathbb{E}[e^{j\omega}]$. Then, (12) can be written as
$\gamma\approx K|{\mathrm{sinc}}(2\pi\epsilon_{\max})|.$ (15)
Furthermore, since $\Lambda_{\max}-\Lambda_{\min}=2\gamma$, we can obtain the estimate of the number of active users, as follows:
$\hat{K}_{\mathrm{prop2}}=\frac{\lambda_{\max}-\lambda_{\min}}{2|{\mathrm{sinc}}(2\pi\epsilon_{\max})|}.$ (16)
This scheme relies solely on the knowledge of $\epsilon_{\max}$, which is usually available as the frequency errors of uplink users are confined within a certain range specified by the standard.
V Numerical Results
In this section, we investigate the performance of the proposed schemes through computer simulations. The simulation setup basically follows TABLE I unless otherwise specified. To assess the accuracy of AUE, we evaluate the normalized estimation error, which is defined as
$E_{K}$ $\triangleq\mathbb{E}\left[\frac{|\hat{K}-K|}{K}\right],$ (17)
where $\hat{K}$ is the estimate of the number of active users $K$. In the simulations, we compare the following methods:
  • Proposed1: the proposed scheme given by (14).
  • Proposed2: the proposed scheme given by (16).
  • Orthogonal: the AUE scheme in [11] using the sequence $\mathbf{s}_{\perp}$ given by (5).
  • MLE: the AUE scheme in [14] based on the MLE given by (7).
The signal-to-noise ratio (SNR) is defined as $1/\sigma^{2}_{\mathrm{z}}$.
Fig. 1 shows the normalized estimation error performance as a function of $\epsilon_{\max}$ for $M=32$ and $128$. As seen in the figure, Proposed1 outperforms the other schemes while Orthogonal performs worse than all the others, especially when $\epsilon_{\max}$ is large. Interestingly, the performance of Proposed1 slightly improves as $\epsilon_{\max}$ increases. This improvement is presumed to be due to the decrease in $\gamma$ caused by the increase in $\epsilon_{\max}$. We can also see that Proposed2 becomes more robust against CFOs than MLE when $M=128$, due to the consideration of the effect of CFOs in the covariance matrix of the received signal.
In Fig. 2, we show the normalized estimation error performance when the number of antennas $M$ varies from 16 to 128. Similar to Fig. 1, this figure demonstrates that Proposed2 outperforms MLE when $M$ is sufficiently large. Moreover, it turns out that Proposed1 consistently achieves better performance than the other schemes.
TABLE I: Simulation parameters.
# of potential users $N$ 100
# of active users $K$ 25
# of antennas at the BS $M$ 32
Maximum normalized CFO $\epsilon_{\max}$ 0.15
SNR $1/\sigma^{2}_{\mathrm{z}}$ 10 [dB]
Figure 1: Estimation performance as a function of $\epsilon_{\max}$.
Figure 2: Estimation performance as a function of $M$.
Next, we examine the effect of the noise variance $\sigma^{2}_{\mathrm{z}}$ in the AUE. Fig. 3 shows the normalized estimation error performance versus SNR. As seen in the figure, the relationship among the performances of all AUE schemes does not vary across all SNR. In addition, Proposed2 follows the same performance trend as the other schemes, even though it does not use the knowledge of $\sigma^{2}_{\mathrm{z}}$.
Finally, Fig. 4 shows the normalized estimation error performance versus the number of active users $K$. This result indicates that the estimation error of the proposed schemes remains roughly constant under different value of $K$, similar to the conventional schemes. In light of the above, we can conclude that the proposed schemes are robust to both CFOs and the number of active users.
Figure 3: Estimation performance as a function of SNR $1/\sigma^{2}_{\mathrm{z}}$.
Figure 4: Estimation performance as a function of $K$.
VI Conclusion
In this letter, we proposed two AUE schemes for GF-NOMA in the presence of CFOs. Both proposed schemes utilize the eigenvalues of the sample covariance matrix of the received signal, with one scheme designed to function without requiring noise variance information. Numerical results demonstrate that the proposed schemes outperform conventional methods when the noise variance is known. Moreover, even without knowledge of the noise variance, our schemes perform comparably to these conventional approaches.

### v2

@@
 The two schemes differ in their treatment of noise variance: one exploits known variance information, while the other is designed to function without this knowledge.
-Simulation results demonstrate the effectiveness of the proposed schemes in terms of the normalized mean-squared error.
+Simulation results demonstrate the effectiveness of the proposed schemes in terms of the normalized root-mean-squared error.
 Index Terms:
@@
 As a promising solution to this problem, grant-free random access schemes such as grant-free non-orthogonal multiple access (GF-NOMA) in which the base station (BS) does not exclusively assign radio resources to active users for data transmission have attracted significant attention [4, 5].
-In a GF-NOMA system, each active user can directly transmit its packet to the BS without waiting for permission; as a result, however, the BS must cope with active user detection (AUD) to recovery transmitted data accurately.
+In a GF-NOMA system, each active user can directly transmit its packet to the BS without waiting for permission; as a result, however, the BS must cope with active user detection (AUD) to recover transmitted data accurately.
 In this context, many approaches to accurate AUD have been investigated for synchronous GF-NOMA systems, where perfect synchronization between the BS and users in both time and frequency is assumed[6, 7, 8, 9, 10, 11, 12, 13, 14].
 On the other hand, achieving ideal synchronization is impractical due to the nature of grant-free transmission, and AUD for asynchronous scenarios remains an ongoing research area[4].
-In practice, due to the practical limitation of the crystal oscillators[15], the resulting carrier frequency offsets (CFOs) among the users degrade the performance.
+In practice, due to practical limitations of crystal oscillators[15], the resulting carrier frequency offsets (CFOs) among the users degrade the performance.
 To overcome this impediment, some approaches to AUD in the presence of CFOs have recently been proposed[16, 17, 18, 19, 20].
@@
 In addition, [17] and [18] handled AUD using maximum likelihood estimation (MLE), which achieves higher accuracy than the BCD-based method of [16].
-Meanwhile, since the MLE-based method use a discrete approximation of the rotation caused by CFOs, the improvement of the performance requires the increase in the discretization size, resulting in high computational complexity.
+Meanwhile, since the MLE-based method uses a discrete approximation of the rotation caused by CFOs, the improvement of the performance requires the increase in the discretization size, resulting in high computational complexity.
 As an alternative, the approximate message passing (AMP)-based approaches for AUD and CE have been proposed in [19] and [20], assuming that the frequency error is limited to a certain range that meets standard requirements.
@@
 These AUE schemes, utilizing a short common pilot, can estimate the number of active users with low complexity.
-To this end, in this letter, we propose two AUE schemes using a short common pilot sequence for estimating the number of active users in the presence of CFOs.
-Specifically, these AUE schemes that use the eigenvalues of the sample covariance matrix of the received signal, with one scheme designed for scenarios where the noise variance information is known, and the other for when it is unknown.
+However, since these AUE schemes use the pilot sequence, their estimation accuracy may degrade when CFO is present, similar to AUD.
+To this end, in this letter, we propose two AUE schemes using a short common pilot sequence to estimate the number of active users in the presence of CFOs.
+Although the proposed schemes use the short common pilot as in [11] and [14], they do not rely on the orthogonality of the sequences or the formulation based on the pure pilot, both of which are affected by CFOs.
+Specifically, the proposed AUE schemes utilize the eigenvalues of the sample covariance matrix of the received signal, with one scheme designed for scenarios where the noise variance information is known, and the other for when it is unknown.
+Furthermore, the computational complexity for these schemes is comparable to that of conventional approaches.
 Numerical results demonstrate that the proposed AUE schemes achieve superior estimation accuracy compared to the conventional schemes proposed in [11] and [14].
-Notation: Throughout this paper, we use the following notation.
-The transpose and conjugate transpose are denoted by $(\cdot)^{\mathrm{T}}$ and $(\cdot)^{\mathrm{H}}$, respectively.
+Notation: The transpose and conjugate transpose are denoted by $(\cdot)^{\mathrm{T}}$ and $(\cdot)^{\mathrm{H}}$, respectively.
 The $l_{p}$-norm of a vector is expressed as $\|\cdot\|_{p}$.
+${}\Re\{\cdot\}$ and $(\cdot)^{\ast}$ denote the real part and conjugate of the argument, respectively.
 In addition, $\mathbf{0}_{L}$ and $\mathbf{I}_{L}$ denote the $L$-dimensional column vectors in which all elements are zeros and the $L\times L$ identity matrix, respectively.
@@
 The expectation operator and operation that rounds $x$ to the nearest integer are represented by $\mathbb{E}[\cdot]$ and $\lfloor x\rceil$, respectively.
+$\mathbf{a}\circ\mathbf{b}$ denotes the element-wise product of two vectors $\mathbf{a}$ and $\mathbf{b}$.
 Finally, $\mathcal{CN}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ represents a multivariate complex Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$.
@@
 In this section, to clarify the difference between conventional and proposed AUE schemes, we review the AUE schemes proposed in [11] and [14].
-These conventional schemes utilize the assumption that the probability function of $\mathbf{Y}$ given $K$ and $\mathbf{s}$ is given by
+These conventional schemes operate under the assumption that the probability function of $\mathbf{Y}$ given $K$, $\mathbf{s}$, and $\omega_{n}=0,\forall n$, is given by
 $p(\mathbf{Y}|K,\mathbf{s})=\frac{1}{|\pi\boldsymbol{\Sigma}|^{M}}\prod^{M}_{m=1}\exp\left(-\mathbf{y}^{\mathrm{H}}_{m}\boldsymbol{\Sigma}^{-1}\mathbf{y}^{\mathrm{H}}_{m}\right),$ (3)
-where
-$\boldsymbol{\Sigma}=\mathbb{E}\left[\mathbf{y}_{m}\mathbf{y}^{\mathrm{H}}_{m}\right]=K\mathbf{s}\mathbf{s}^{\mathrm{H}}+\sigma^{2}_{\mathrm{z}}\mathbf{I}_{2}.$ (4)
-Note that this assumption does not consider the effect of CFOs.
+where $\boldsymbol{\Sigma}=\mathbb{E}\left[\mathbf{y}_{m}\mathbf{y}^{\mathrm{H}}_{m}\right]=K\mathbf{s}\mathbf{s}^{\mathrm{H}}+\sigma^{2}_{\mathrm{z}}\mathbf{I}_{2}.$ Thus, this assumption does not consider the effect of CFOs.
 The AUE scheme in [11] utilizes the sequence $\mathbf{s}_{\perp}\in\mathbb{C}^{2\times 1}$ that is orthogonal to $\mathbf{s}$ and satisfies $\|\mathbf{s}_{\perp}\|_{2}=\|\mathbf{s}\|_{2}$.
@@
 $\hat{K}_{\mathrm{orth}}=\Bigg{\lfloor}\frac{\mathbf{s}^{\mathrm{H}}\mathbf{R}\mathbf{s}}{\|\mathbf{s}\|^{4}_{2}}-\frac{\mathbf{s}^{\mathrm{H}}_{\perp}\mathbf{R}\mathbf{s}_{\perp}}{\|\mathbf{s}\|^{2}_{2}\|\mathbf{s}_{\perp}\|^{2}_{2}}\Bigg{\rceil}.$ (4)
-When we set $\|\mathbf{s}\|_{2}=\|\mathbf{s}_{\perp}\|_{2}=1$, (5) is equivalent to [11, eq.
+When $\|\mathbf{s}\|_{2}=\|\mathbf{s}_{\perp}\|_{2}=1$, (4) is equivalent to [11, eq.
 (5)].
-It is worth noting that this scheme does not require the knowledge of noise variance and that $\hat{K}_{\mathrm{orth}}$ converges to $K$ when $M\to\infty$ and $\omega_{n}=0$.
+It is worth noting that this scheme does not require the knowledge of noise variance $\sigma^{2}_{\mathrm{z}}$.
 On the other hand, the AUE scheme in [14] considers the MLE problem of $K$. This problem can be formulated as
@@
 $\hat{K}_{\mathrm{MLE}}=\Bigg{\lfloor}\frac{\mathbf{s}^{\mathrm{H}}\mathbf{R}\mathbf{s}}{\|\mathbf{s}\|^{4}_{2}}-\frac{\sigma^{2}_{\mathrm{z}}}{\|\mathbf{s}\|^{2}_{2}}\Bigg{\rceil}.$ (6)
-As is obvious from (7), the estimation of $K$ in [14] does not rely on the sequence orthogonal to $\mathbf{s}$, while it requires the knowledge of noise variance.
+As is obvious from (6), the estimation of $K$ in [14] does not rely on $\mathbf{s}_{\perp}$, while it requires $\sigma^{2}_{\mathrm{z}}$.
 IV Proposed AUE Schemes
-In this section, we propose two AUE schemes using the sample covariance matrix of the received signal.
-Specifically, the proposed schemes utilize the eigenvalues of $\mathbf{R}$.
-Hereafter, $\lambda_{\max}$ and $\lambda_{\min}\ (<\lambda_{\max})$ denote the eigenvalues.
+In this section, we propose two AUE schemes, named Eig-sum and Eig-diff, which utilize the eigenvalues of the sample covariance matrix $\mathbf{R}$, $\lambda_{\max}$ and $\lambda_{\min}\ (<\lambda_{\max})$.
 IV-A Preliminaries
@@
 $\approx K\sqrt{\mathbb{E}\left[e^{j\omega}\right]\mathbb{E}\left[e^{-j\omega}\right]},$ (11)
-where $\omega$ represents the random variable following a uniform distribution in the interval $[-2\pi\epsilon_{\max},2\pi\epsilon_{\max}]$.
-$\mathbb{E}\left[e^{j\omega}\right]$ in (12) is the characteristic function of $\omega$, which is given by
+where $\omega$ represents a random variable that takes values within the interval $[-2\pi\epsilon_{\max},2\pi\epsilon_{\max}]$.
+$\mathbb{E}\left[e^{j\omega}\right]$ in (11) is the characteristic function of $\omega$.
+IV-B Eig-sum: Proposed AUE With Knowledge of $\sigma^{2}_{\mathrm{z}}$
+As is obvious from (8) and (9), $K$ can ideally be obtained based on the sum of the eigenvalues of $\tilde{\boldsymbol{\Sigma}}$, i.e., $K=(\Lambda_{\max}+\Lambda_{\min})/2-\sigma^{2}_{\mathrm{z}}$, if the information of $\sigma^{2}_{\mathrm{z}}$ is available at the BS.
+Following this fact, Eig-sum obtains the estimate of $K$ by
+${\color[rgb]{0,0,0}\hat{K}_{\mathrm{sum}}=\Bigg{\lfloor}\frac{\lambda_{\max}+\lambda_{\min}}{2}-\sigma^{2}_{\mathrm{z}}\Bigg{\rceil}.}$ (12)
+Since this scheme is based on the calculation to offset the term including CFOs, $\gamma$, it is expected to be robust against CFOs.
+As in [11], we consider the normalized root mean-squared error (NRMSE) of this scheme, which is defined as
+${\color[rgb]{0,0,0}\mathrm{NRMSE}=\frac{1}{K}\sqrt{\mathbb{E}\left[\left(\hat{K}-K\right)^{2}\right]},}$ (13)
+where $\hat{K}$ is the estimate of the number of active users $K$. Given $\alpha=\mathbb{E}[e^{j\omega}]$, the NRMSE of Eig-sum can be obtained by
+$\mathrm{NRMSE}=\frac{1}{K}\sqrt{\frac{K+K(K-1)\alpha^{2}+\left(K+\sigma^{2}_{\mathrm{z}}\right)^{2}}{2M}}.$ (14)
+For the derivation of (14), please refer to Appendix.
+IV-C Eig-diff: Proposed AUE Without Knowledge of $\sigma^{2}_{\mathrm{z}}$
+Since $\mathbb{E}[e^{j\omega}]$ is an even function, $\mathbb{E}[e^{-j\omega}]=\mathbb{E}[e^{j\omega}]$ is satisfied.
+[^1] In the literature, CFO is often modeled as a random variable that follows a symmetric distribution, such as a uniform or Gaussian distribution.
+Hence, the characteristic function of $\omega$ is real-valued and even.
+Then, (11) can be written as
+${\color[rgb]{0,0,0}\gamma\approx K|\mathbb{E}[e^{j\omega}]|.}$ (15)
+Furthermore, since $\Lambda_{\max}-\Lambda_{\min}=2\gamma$, we can obtain the estimate of the number of active users, as follows:
+${\color[rgb]{0,0,0}\hat{K}_{\mathrm{diff}}=\Bigg{\lfloor}\frac{\lambda_{\max}-\lambda_{\min}}{2|\mathbb{E}[e^{j\omega}]|}\Bigg{\rceil}.}$ (16)
+This scheme relies solely on the knowledge of $\epsilon_{\max}$, which is usually available as the frequency errors of uplink users are confined within a certain range specified by the standard.
+Note that, since the analysis of the NRMSE of Eig-diff is very complicated, this is left for future work.
+IV-D Complexity Comparison
+In this subsection, we discuss the computational complexity required for the proposed and conventional AUE schemes. Here, we express the entries of $\mathbf{R}$, as follows:
 […]
-$=\frac{\sin(2\pi\epsilon_{\max})}{2\pi\epsilon_{\max}}$
-$=\mathrm{sinc}(2\pi\epsilon_{\max}).$ (13)
-IV-B Proposed AUE Scheme With Knowledge of $\sigma^{2}_{\mathrm{z}}$
-As is obvious from (9) and (10), $K$ can be ideally obtained based on the sum of eigenvalues of $\tilde{\boldsymbol{\Sigma}}$, i.e., $K=(\Lambda_{\max}+\Lambda_{\min})/2-\sigma^{2}_{\mathrm{z}}$, if the information of $\sigma^{2}_{\mathrm{z}}$ is available at the BS.
-Following this fact, one of the proposed schemes obtains the estimate of $K$ by
-$\hat{K}_{\mathrm{prop1}}=\frac{\lambda_{\max}+\lambda_{\min}}{2}-\sigma^{2}_{\mathrm{z}}.$ (14)
-Since this scheme is based on the calculation to offset the term including CFOs, $\gamma$, it is expected to be robust against CFOs.
-IV-C Proposed AUE Scheme Without Knowledge of $\sigma^{2}_{\mathrm{z}}$
-Since $\mathbb{E}[e^{j\omega}]$ is an even function, $\mathbb{E}[e^{-j\omega}]=\mathbb{E}[e^{j\omega}]$. Then, (12) can be written as
-$\gamma\approx K|{\mathrm{sinc}}(2\pi\epsilon_{\max})|.$ (15)
-Furthermore, since $\Lambda_{\max}-\Lambda_{\min}=2\gamma$, we can obtain the estimate of the number of active users, as follows:
-$\hat{K}_{\mathrm{prop2}}=\frac{\lambda_{\max}-\lambda_{\min}}{2|{\mathrm{sinc}}(2\pi\epsilon_{\max})|}.$ (16)
-This scheme relies solely on the knowledge of $\epsilon_{\max}$, which is usually available as the frequency errors of uplink users are confined within a certain range specified by the standard.
+where $\mathbf{y}_{1}\in\mathbb{C}^{1\times M}$ and $\mathbf{y}_{2}\in\mathbb{C}^{1\times M}$ represent the first and second rows of the received signal $\mathbf{Y}$, respectively.
+Then, $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ are given by
+$\mathbf{y}_{1}=\sum_{n\in\mathcal{A}}\mathbf{h}^{\mathrm{T}}_{n}+\mathbf{z}_{1},\quad\mathbf{y}_{2}=\sum_{n\in\mathcal{A}}e^{j\omega_{n}}\mathbf{h}^{\mathrm{T}}_{n}+\mathbf{z}_{2},$ (18)
+where $\mathbf{z}_{1}\in\mathbb{C}^{1\times M}$ and $\mathbf{z}_{2}\in\mathbb{C}^{1\times M}$ are the first and second rows of the noise matrix $\mathbf{Z}$.
+[…]
+TABLE I: Computational Complexity of AUE Schemes.
+AUE scheme Number of multiplications
+Eig-sum: $\left\lfloor\frac{R_{1}+R_{2}}{2}-\sigma^{2}_{\mathrm{z}}\right\rceil$ $32M+$ 3
+Eig-diff: $\left\lfloor\frac{\sqrt{(R_{1}-R_{2})^{2}+4\tilde{R}\tilde{R}^{\ast}}}{2\mathbb{E}[e^{j\omega}]}\right\rceil$ $73M+$ 7
+Orthogonal[11]: $\left\lfloor\Re\{\tilde{R}\}\right\rceil$ $M+1$
+MLE[14]: $\left\lfloor\frac{R_{1}+R_{2}+2\Re\{\tilde{R}\}}{4}-\frac{\sigma^{2}_{\mathrm{z}}}{2}\right\rceil$ $53M+$ 5
 V Numerical Results
+TABLE II: Simulation Parameters.
+Number of potential users $N$ 100
+Number of active users $K$ 25
+Number of antennas at the BS $M$ 32
+Maximum normalized CFO $\epsilon_{\max}$ 0.15
+Distribution of $\omega_{n}$ Uniform distribution:
+$\omega_{n}\in[-2\pi\epsilon_{\max},2\pi\epsilon_{\max}]$
+SNR $1/\sigma^{2}_{\mathrm{z}}$ 10 [dB]
 In this section, we investigate the performance of the proposed schemes through computer simulations.
-The simulation setup basically follows TABLE I unless otherwise specified.
-To assess the accuracy of AUE, we evaluate the normalized estimation error, which is defined as
-$E_{K}$ $\triangleq\mathbb{E}\left[\frac{|\hat{K}-K|}{K}\right],$ (17)
-where $\hat{K}$ is the estimate of the number of active users $K$. In the simulations, we compare the following methods:
-  • Proposed1: the proposed scheme given by (14).
-  • Proposed2: the proposed scheme given by (16).
-  • Orthogonal: the AUE scheme in [11] using the sequence $\mathbf{s}_{\perp}$ given by (5).
-  • MLE: the AUE scheme in [14] based on the MLE given by (7).
-The signal-to-noise ratio (SNR) is defined as $1/\sigma^{2}_{\mathrm{z}}$.
+The simulation setup basically follows TABLE II unless otherwise specified.
+To assess the accuracy of AUE, we evaluate the NRMSE performance, which is defined by (13).
+In the simulations, we compare the following methods:
+  • Eig-sum: the proposed scheme given by (12).
+  • Eig-diff: the proposed scheme given by (16).
+  • Orthogonal: the AUE scheme in [11] given by (4).
+  • MLE: the AUE scheme in [14] given by (6).
+The signal-to-noise ratio (SNR) is defined as $1/\sigma^{2}_{\mathrm{z}}$. In addition, all simulation results include the analyzed NRMSE of Eig-sum in (14) as “Eig-sum (theory).”
 Fig.
-1 shows the normalized estimation error performance as a function of $\epsilon_{\max}$ for $M=32$ and $128$.
-As seen in the figure, Proposed1 outperforms the other schemes while Orthogonal performs worse than all the others, especially when $\epsilon_{\max}$ is large.
-Interestingly, the performance of Proposed1 slightly improves as $\epsilon_{\max}$ increases.
+1 shows the NRMSE performance as a function of $\epsilon_{\max}$ for $M=32$ and $M=128$.
+Figs.
+1(a) and 1(b) show the performance for uniform CFO and Gaussian CFO  [^2] In this case, $\omega_{n}$ follows the Gaussian distribution with mean zero and standard deviation $32\pi\epsilon_{\max}/$ 3, in which about 99.7% of the values lie within $[-2\pi\epsilon_{\max},2\pi\epsilon_{\max}]$.
+cases, respectively.
+As seen in Figs.
+1(a) and 1(b), Eig-sum outperforms the other schemes, while Orthogonal performs worse than all the others, especially when $\epsilon_{\max}$ is large.
+Interestingly, the performance of Eig-sum slightly improves as $\epsilon_{\max}$ increases.
 This improvement is presumed to be due to the decrease in $\gamma$ caused by the increase in $\epsilon_{\max}$.
-We can also see that Proposed2 becomes more robust against CFOs than MLE when $M=128$, due to the consideration of the effect of CFOs in the covariance matrix of the received signal.
+Furthermore, the analyzed NRMSE aligns with the simulated NRMSE.
+We can also see that the performance of Eig-diff is superior to or comparable to that of MLE when $M=128$, despite not using $\sigma^{2}_{\mathrm{z}}$.
+This is because Eig-diff considers the effect of CFOs in the covariance matrix of the received signal, making it more robust.
 In Fig.
-2, we show the normalized estimation error performance when the number of antennas $M$ varies from 16 to 128.
+4, we show the NRMSE performance when the number of antennas $M$ varies from 16 to 128.
 Similar to Fig.
-1, this figure demonstrates that Proposed2 outperforms MLE when $M$ is sufficiently large.
-Moreover, it turns out that Proposed1 consistently achieves better performance than the other schemes.
-TABLE I: Simulation parameters.
-# of potential users $N$ 100
-# of active users $K$ 25
-# of antennas at the BS $M$ 32
-Maximum normalized CFO $\epsilon_{\max}$ 0.15
-SNR $1/\sigma^{2}_{\mathrm{z}}$ 10 [dB]
-Figure 1: Estimation performance as a function of $\epsilon_{\max}$.
-Figure 2: Estimation performance as a function of $M$.
+1, this figure demonstrates that Eig-diff performs comparably to MLE when $M$ is sufficiently large.
+Moreover, it turns out that Eig-sum consistently outperforms the other schemes.
+(a) Uniform CFO
+(b) Gaussian CFO
+Figure 1: NRMSE performance versus $\epsilon_{\max}$.
 Next, we examine the effect of the noise variance $\sigma^{2}_{\mathrm{z}}$ in the AUE.
 Fig.
-3 shows the normalized estimation error performance versus SNR.
+4 shows the NRMSE performance versus SNR.
 As seen in the figure, the relationship among the performances of all AUE schemes does not vary across all SNR.
-In addition, Proposed2 follows the same performance trend as the other schemes, even though it does not use the knowledge of $\sigma^{2}_{\mathrm{z}}$.
+Given that the NRMSE of Eig-sum remains nearly constant with changes in SNR, especially in the high SNR region, it is reasonable to presume that the NRMSE of other methods would similarly remain largely unaffected by SNR variations.
 Finally, Fig.
-4 shows the normalized estimation error performance versus the number of active users $K$.
+4 shows the NRMSE performance versus the number of active users $K$.
 This result indicates that the estimation error of the proposed schemes remains roughly constant under different value of $K$, similar to the conventional schemes.
 In light of the above, we can conclude that the proposed schemes are robust to both CFOs and the number of active users.
-Figure 3: Estimation performance as a function of SNR $1/\sigma^{2}_{\mathrm{z}}$.
-Figure 4: Estimation performance as a function of $K$.
+Figure 2: NRMSE performance versus $M$.
+Figure 3: NRMSE performance versus SNR.
+Figure 4: NRMSE performance versus $K$.
 VI Conclusion
@@
 Moreover, even without knowledge of the noise variance, our schemes perform comparably to these conventional approaches.
+Although this study focuses on a scenario with CFOs, addressing timing offsets is also necessary and is therefore left for future work.
+The NRMSE of Eig-sum can be calculated by (19), which is based on $\hat{K}_{\mathrm{sum}}=\lfloor(R_{1}+R_{2})/2-\sigma^{2}_{\mathrm{z}}\rceil$.
+${\color[rgb]{0,0,0}\mathrm{NRMSE}=\frac{1}{K}\sqrt{\mathbb{E}\left[\left(\frac{R_{1}+R_{2}}{2}-\sigma^{2}_{\mathrm{z}}\right)^{2}\right]-K^{2}}=\frac{1}{K}\sqrt{\frac{1}{4}\mathbb{E}\left[R^{2}_{1}\right]+\frac{1}{4}\mathbb{E}\left[R^{2}_{2}\right]+\frac{1}{2}\mathbb{E}\left[R_{1}R_{2}\right]-\sigma^{2}_{\mathrm{z}}\left(\mathbb{E}\left[R_{1}\right]+\mathbb{E}\left[R_{2}\right]\right)+\sigma^{4}_{\mathrm{z}}-K^{2}}.}$ (19)
+Note that the derivation of (14) ignores the rounding operation as in [11].
+We first consider $\mathbb{E}[R_{1}]$, $\mathbb{E}[R_{1}]$, $\mathbb{E}[R^{2}_{1}]$, and $\mathbb{E}[R^{2}_{2}]$.
+Since $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ can be regarded as a random vector that follows $\mathcal{CN}(\mathbf{0}_{M},(K+\sigma^{2}_{\mathrm{z}})\mathbf{I}_{M})$, $R_{1}$ and $R_{2}$ follow the Erlang distribution with the shape parameter $M$ and rate parameter $M/(K+\sigma^{2}_{\mathrm{z}})$.
+The mean and variance of this distribution are $K+\sigma^{2}_{\mathrm{z}}$ and $(K+\sigma^{2}_{\mathrm{z}})^{2}/M$, respectively.
+Consequently, we can obtain the following relationships.
+$\mathbb{E}[R_{1}]$ $=\mathbb{E}[R_{2}]=K+\sigma^{2}_{\mathrm{z}},$ (20)
+$\mathbb{E}[R^{2}_{1}]$ $=\mathbb{E}[R^{2}_{2}]=\left(1+\frac{1}{M}\right)\left(K+\sigma^{2}_{\mathrm{z}}\right)^{2}.$ (21)
+Next, we discuss the value of $\mathbb{E}[R_{1}R_{2}]$. $R_{1}$ and $R_{2}$ can be expanded as follows:
+$R_{1}$ $=\sum_{n\in\mathcal{A}}\frac{\|\mathbf{h}_{n}\|^{2}_{2}}{M}+\sum_{n\in\mathcal{A}}\sum_{n^{\prime}\in\mathcal{A},n^{\prime}\neq n}\frac{\mathbf{h}^{\mathrm{T}}_{n}\mathbf{h}^{\ast}_{n^{\prime}}}{M}$
+$\quad+\sum_{n\in\mathcal{A}}\frac{\mathbf{h}^{\mathrm{T}}_{n}\mathbf{z}^{\mathrm{H}}_{1}}{M}+\sum_{n\in\mathcal{A}}\frac{\mathbf{z}_{1}\mathbf{h}^{\ast}_{n}}{M}+\frac{\|\mathbf{z}_{1}\|^{2}_{2}}{M},$ (22)
+$R_{2}$ $=\sum_{n\in\mathcal{A}}\frac{\|\mathbf{h}_{n}\|^{2}_{2}}{M}+\sum_{n\in\mathcal{A}}\sum_{n^{\prime}\in\mathcal{A},n^{\prime}\neq n}e^{j(\omega_{n}-\omega_{n^{\prime}})}\frac{\mathbf{h}^{\mathrm{T}}_{n}\mathbf{h}^{\ast}_{n^{\prime}}}{M}$
+$\quad+\sum_{n\in\mathcal{A}}e^{j\omega_{n}}\frac{\mathbf{h}^{\mathrm{T}}_{n}\mathbf{z}^{\mathrm{H}}_{2}}{M}+\sum_{n\in\mathcal{A}}e^{-j\omega_{n}}\frac{\mathbf{z}_{2}\mathbf{h}^{\ast}_{n}}{M}+\frac{\|\mathbf{z}_{2}\|^{2}_{2}}{M}.$ (23)
+Given that $\mathbf{h}_{n}$, $\omega_{n}$, $\mathbf{z}_{1}$, and $\mathbf{z}_{2}$ are independent of each other, $R_{1}R_{2}$ can be expressed by (24) without including negligible terms,
+${\color[rgb]{0,0,0}R_{1}R_{2}=\!\sum_{n\in\mathcal{A}}\frac{\|\mathbf{h}_{n}\|^{4}_{2}}{M^{2}}+\!\sum_{n\in\mathcal{A}}\sum_{n^{\prime}\in\mathcal{A},n^{\prime}\neq n}\left(\frac{\|\mathbf{h}_{n}\|^{2}_{2}\|\mathbf{h}_{n^{\prime}}\|^{2}_{2}}{M^{2}}+e^{j(\omega_{n}-\omega_{n^{\prime}})}\frac{\mathbf{h}^{\mathrm{T}}_{n^{\prime}}\mathbf{h}^{\ast}_{n^{\prime}}\mathbf{h}^{\mathrm{T}}_{n^{\prime}}\mathbf{h}^{\ast}_{n}}{M^{2}}\right)+\frac{\|\mathbf{z}_{1}\|^{2}_{2}+\|\mathbf{z}_{2}\|^{2}_{2}}{M^{2}}\!\sum_{n\in\mathcal{A}}\|\mathbf{h}_{n}\|^{2}_{2}+\frac{\|\mathbf{z}_{1}\|^{2}_{2}\|\mathbf{z}_{2}\|^{2}_{2}}{M^{2}}.}$ (24)
+where we use the following relationship:
+$\sum_{n\in\mathcal{A}}\sum_{n^{\prime}\in\mathcal{A},n^{\prime}\neq n}\frac{\mathbf{h}^{\mathrm{T}}_{n}\mathbf{h}^{\ast}_{n^{\prime}}}{M}=\sum_{n\in\mathcal{A}}\sum_{n^{\prime}\in\mathcal{A},n^{\prime}\neq n}\frac{\mathbf{h}^{\mathrm{T}}_{n^{\prime}}\mathbf{h}^{\ast}_{n}}{M}.$ (25)
+In addition, since $\mathbf{h}_{n}\sim\mathcal{CN}(\mathbf{0}_{M},\mathbf{I}_{M})$, $\mathbf{z}_{1}\sim\mathcal{CN}(\mathbf{0}_{M},\sigma^{2}_{\mathrm{z}}\mathbf{I}_{M})$, and $\mathbf{z}_{2}\sim\mathcal{CN}(\mathbf{0}_{M},\sigma^{2}_{\mathrm{z}}\mathbf{I}_{M})$, we can obtain the relationships: $\mathbb{E}\left[\|\mathbf{h}_{n}\|^{4}_{2}\right]=M+M^{2}$, $\mathbb{E}\left[\|\mathbf{h}_{n}\|^{2}_{2}\right]=M$, $\mathbb{E}\left[\mathbf{h}^{\ast}_{n^{\prime}}\mathbf{h}^{\mathrm{T}}_{n^{\prime}}\right]=\mathbf{I}_{M}$, and $\mathbb{E}\left[\|\mathbf{z}_{1}\|^{2}_{2}\right]=\mathbb{E}\left[\|\mathbf{z}_{2}\|^{2}_{2}\right]=M\sigma^{2}_{\mathrm{z}}$.
+As a result, $\mathbb{E}[R_{1}R_{2}]$ can be calculated by
+$\mathbb{E}\left[R_{1}R_{2}\right]=\frac{K+K(K-1)\alpha^{2}}{M}+\left(K+\sigma^{2}_{\mathrm{z}}\right)^{2}.$ (26)
+In light of the above, by substituting (20), (21), and (26) into (19), we can derive (14).